(A) How much energy is released when 0.500g of B-11 and an excess of H-1 are fused to produce three alpha particles? The masses are: B-11 = 11.009305amu, H-1 = 1.007825amu, and He-4 = 4.002603.
Answer #1

Δm = 11.009305 + 1.007825amu - 3*4.002603
= 9.321*10^-3 amu

1 amu = 931.5 MeV
9.321*10^-3 amu = 9.321*10^-3 amu * 931.5 = 8.68 MeV

Answer: 8.68 MeV